(2) Consider the following game of chance. You pay $30 to enter the game. First, you choose a number at random from {1, 2, ..(c) What is the variance and standard deviation of your gain in this game?
Answer #1

L 10 | - 90x21 - 2 20 = E(XIE(y)-So I independent I = 20+ (90+1)(a & Game fee = $30 - Random No from {1, 2, -,204 P(x:x1 = 1